OpenStudy (anonymous):

An item that originally cost $100 is decreased by 8%. The reduced price is then increased by 8%. The resulting price is _____. $100 $92 between $92 and $100 more than $100

OpenStudy (whpalmer4):

Reducing price by 8% is the same as multiplying by (100-8)/100. Increasing price by 8% is the same as multiplying by (100+8)/100. So, we start out with $100, decrease by 8%, increase by 8%, we get: 100 * (100-8)/100*(100+8)/100 = (100*92*108)/(100*100) = 99.36 This is why it is much easier to do well in the stock market if you avoid losing money! If you lose 8% of your money, to get back to even you need to earn 8.7%. If you lose half of your money (50%), you need to earn a whopping 200% to get back to where you started!